OpenStudy (anonymous):

Simplify. r^2s^5t^4 / r^2s^2t^5 A. s^3/t^2 B. s^3/t C. s^2/t^3 D. s/t

OpenStudy (zzr0ck3r):

\(\large \frac{r^2s^5t^4}{r^2s^2t^5}=\frac{r^2}{r^2}\frac{s^5}{s^2}\frac{t^4}{t^5}\) now \(\frac{a^b}{a^c}=a^{b-c}=\frac{1}{a^{c-b}}\)

OpenStudy (zzr0ck3r):

so @eflatt13 what is \(\large \frac{r^2}{r^2}= \ ?\)

OpenStudy (anonymous):

s^3/t?